5/22: RT PC maintenance -> go to safe mode for all VIS connected Dolphin (BS)
5/22-23: MZ release from PSL
5/20-24: Exchanging cryo valves.
- 5/27-6/2: Y arm mass lock.
- 6/3-6/14: Installation of TMS for ETMY
- 6/8: Engineering run
5/22AM~: tuned off cryostat cooler, finesse measurement for Xarm
5/22AM~: tuned off shield cooler, finesse measurement for Xarm
5/2: Cooling start
- 5/29: Cooling done
5/23AM~: tuned off cryostat cooler, finesse measurement for Yarm
5/17~: tuned off ETMY cryostat cooler
5/24PM~ inject 100Pa Nitrogen. -> postponed after 6/3?
5/27 evacuation of Nitrogen. -> postponed after 6/3?
- Long term:
- 4/22-5/31: Yarm commissioning
- 5/27-6/30: DRMI commissioning
- 7/1-8/9: DRFPMI commissioning
- 8/12-10/4: DRFPMI ASC
- 10/7-11/15: Post Commissioning
- 11/18-12/13: Noise hunting
- IOO maybe
Report of work
- COM: Finesse measurement did not succeed by some troubles.
- T cam fo X arm had trouble. It was fixed this morning by restarting Team.
- SR3 drift caused by VIS work in the mid night. Working time should be coordinated.
- Type A: After stopping cryo cooler of ITMY stopped and temperature went up, ITMY surface looks better by T cam image with green light. Spots by dust? seems to be gone and the surface looks more clean.
- IOO: Alignment of PSL is back after MZ was removed day before yesterday.
- DGS: Maintenance work afternoon for k1asc0.